Mary is 5 years younger than alison....that total of the 2 girls age is 15.Sherry is 15 years older than Alison,how old is sherry
Answer by ankor@dixie-net.com(22740)

Mary is 5 years younger than alison..
m = a - 5
or
m - a = -5
..that total of the 2 girls age is 15.
m + a = 15
m - a = -5
--------------------Adding eliminates a, find m
2m + 0 = 10
m = 10/2
m = 5 yrs is Mary's age
then "Mary is 5 years younger than alison.." therefore:
a = 5 + 5
a = 10 yrs is Alison's age
:
Sherry is 15 years older than Alison,
s = a + 15
s = 10 + 15
s = 25 yrs is Sherry's age